OpenStudy (anonymous):

Solve (6x + 30) = x + 5(x + 4) - 2x my options are A. 6 B. 0 C. All Numbers D. No Solution

OpenStudy (anonymous):

remove the parentheses first using the distributive law

OpenStudy (anonymous):

then combine like terms let me know what you get then you can starting adding and subtracting from both sides

OpenStudy (anonymous):

6x+30= x+5x+20-2x 6x-x+2x=20-30 7x=-10 x=-10/7

OpenStudy (anonymous):

some mistake there, try it yourself

pooja195 (pooja195):

Hint: Mistake is in line 2

OpenStudy (anonymous):

I got x = -5

OpenStudy (anonymous):

But my options are A. 6 B. 0 C. All Numbers D. No Solution

OpenStudy (dayakar):

6x+30 =x+5x+20-2x 6x+30=4x+20

OpenStudy (dayakar):

6x-4x=20-30

OpenStudy (dayakar):

2x=-10

OpenStudy (dayakar):

x= -10/2 x=-5